Make-or-Buy Decision

Zion Manufacturing had always made its components in-house. However, Bryce Component Works had recently offered to supply one component, K2, at a price of $12 each. Zion uses 4,900 units of Component K2 each year. The cost per unit of this component is as follows:

Direct materials $7.65
Direct labor 2.80
Variable overhead 1.40
Fixed overhead 3.00
   Total $14.85

The fixed overhead is an allocated expense; none of it would be eliminated if production of Component K2 stopped.

Required:

1. What are the alternatives facing Zion Manufacturing with respect to production of Component K2?

2. List the relevant costs for each alternative. If required, round your answers to the nearest cent.

Total Relevant Cost
Make $ per unit
Buy $ per unit
Differential Cost to Make $ per unit

If Zion decides to purchase the component from Bryce, by how much will operating income increase or decrease?
$

3. Conceptual Connection: Which alternative is better?
Answer #1

1. What are the alternatives facing Zion Manufacturing with respect to production of Component K2?

The alternatives that Zion Manufacturing is facing with respect to production of Component K2is either to make the component inhouse or to buy the component from Bryce.

2. List the relevant costs for each alternative.

Make - DM 7.65 + DL 2.8 + VOH 1.4 = $ 11.85

Buy - $ 12

Differential Cost to Make ( 11.85 - 12 ) = $ 0.15

If Zion decides to purchase the component from Bryce, by how much will operating income increase or decrease?
operating income will decrease by ( 4900 * 0.15 ) = $ 735

3. Conceptual Connection: Which alternative is better?

Making inhouse